MAT240 – Algebra I[36L/24T]  2017

MAT240 - Algebra I[36L/24T]

Instructor

Instructor  Eckhard Meinrenken
Email  mein@math.toronto.edu
Office  BA 6112
Office Hours Thursday 1:30-2:30, or by appointment

Teaching Assistants

Instructor  Debanjana Kundu
Email  dkundu@math.toronto.edu
Office Hours Thursdays 3-4 HU1012 (GANITA lab)

Instructor  Jeff Pike
Email  jpike@math.toronto.edu
Office Hours Tuesdays, 3:30-4:30 Math Aid Centre
Instructor  Ivan Telpukhovsky
Email  ivantelp@math.toronto.edu
Office Hours Monday 4-5PM Math Aid Center

Course Description

Abstract (from academic calendar): A theoretical approach to: vector spaces over arbitrary fields, including ℂ and ℤ_{p}. Subspaces, bases and dimension. Linear transformations, matrices, change of basis, similarity, determinants. Polynomials over a field (including unique factorization, resultants). Eigenvalues, eigenvectors, characteristic polynomial, diagonalization. Minimal polynomial, Cayley-Hamilton theorem.

Textbook. Linear Algebra (4th edition); by Friedberg, Insel and Spence. Prentice Hall

Course Website. The website for the course is

           http://www.math.toronto.edu/mein/teaching/MAT240/MAT240.html

MAT240 – Algebra I[36L/24T]  2017

Homework Assignments

There will be 9 homework assignments, which you must submit by the due date. Your homework mark will be determined by the taking the average of the best 8 assignments. We intend to use the Crowdmark system for the submission of homeworks. For each assignment, you will receive a personal link through which you can submit your solution. Do not forward your link to others. It is important that you submit solutions on time, and that your solutions are legible (use the preview option to verify). If you encounter problems with the crowdmark submission, please let me know by email. While you can certainly discuss homework with classmates, you have to write up the solutions yourself, in your own words. Otherwise it is considered unauthorized aid or assistance (working too closely with another student on an individual assignment so that the end result is too similar), which is an academic offence under the University’s Code of Behaviour on Academic Matters. If you find the solutions in books or on the internet, you must quote your source (and still write it up in your own words!) There will be some additional assignments that are not to be submitted. You are expected to work on these assignments as well. (They may be relevant for the term test or final exam.)

Term Tests

There will be 1 term test, taking place on Tuesday, October 17, 2017, 11:10-13:00 in EX 100. You must bring your student card to the term test. No aids will be allowed. In particular, electronic gadgets (e.g. calculators, phones, smart watches, etc) are not allowed.

Missing Term Work

If you cannot show up for a test because of illness or any other special reason, you MUST submit the official UofT medical certificate, which can be downloaded at

     http://www.illnessverification.utoronto.ca/

There will be NO make-up tests. The marking scheme will be adjusted appropriately for students who have missed a test because of illness or any other (approved) legitimate reason.

MAT240 – Algebra I[36L/24T]  2017

Final Exam

The final exam will take place during the examination period, and will be 3h long. It will cover all the material presented in lectures and tutorials. Aids permitted: None

Marking Scheme

Your final grade is determined in the following way:

Homework  (best 8 out of 9)  20 %
Term Test 1  10/17/2017  30 %
Final Exam  Date: TBA  50 %

100 %

Code of Behaviour / Plagiarism

Students should become familiar with and are expected to adhere to the Code of Behaviour on Academic Matters which can be found at:

     http://www.governingcouncil.utoronto.ca/policies/behaveac.htm